Find the value of k if 5x +7y = k and 10x + 14y = 3 have infinite number of solutions.
Answer by stanbon(75887)

Find the value of k if
5x +7y = k
10x + 14y = 3
------
Modify for elimination::
10x + 14y = 2k
10x + 14y = 3
-----
Subtract to get:
0 = 2k-3
k = 3/2
=============
Cheers,
Stan H.
------------